It is claimed that the average time it takes high school seniors to complete a standardized

test is more than 30 minutes. A random sample of 24 high school seniors took an average

of 30 minutes to complete this test with a standard deviation of 3 minutes. Test the claim

at 1% level of significance.



1
Expert's answer
2021-11-30T11:38:05-0500

Null hypothesis "H_0:\mu=30."

Alternative hypothesis "H_a:\mu>30."

Test statistic: "t=\frac{\bar x-\mu}{\frac{s}{\sqrt{n}}}=\frac{30-30}{\frac{3}{\sqrt{24}}}=0."

Degrees of freedom: "df=n-1=24-1=23."

P-value: "p=P(T>0)=0.5."

Since the p-value is greater than 0.01, fail to reject the null hypothesis.

There is no sufficient evidence that the average time it takes high school seniors to complete a standardized test is more than 30 minutes.